Let X1, . . . , Xn be a random sample from a triangular probability distribution whose density function and moments are:

fX(x) = \tiny \frac{2(b-x)}{b^2} * I{0 \tiny \leq x \tiny \leq b}

a. Find the mean µ of this probability distribution.  

b. Find the Method Of Moments estimator µ(hat) of µ.

c. Is µ(hat) unbiased?

d. Find the Median of this probability distribution.

I will thumbs up any portion or details of how to do this problem, thanks!
Answer #1

fra 2 (box) osasb 6 E(X)= x fcandx Sa 06-3°)dx 6 method of moment for all equaliting sample moment to population moment. Pe l//= cusx) d xockets (2bm.com ] a bh.me 2m a b²+2m² 26² & umb=b²+2m2 umbe-b²=0m² bum-b-am? 2m²- umb+b² = 0 m-20mb + b² = 0 3 tHii dear I will give my 100% can you please like it.??